OpenStudy (anonymous):

Marv is considering two different vehicles. The Cheetah sells for $12,000 and depreciates at a rate of 18% per year. The Panther sells for $18,000 and depreciates at a rate of 24% per year. Which car is worth more after 6 years? Answer Panther, at $3469 Cheetah, at $3469 Panther, at $3648 Cheetah, at $3648

OpenStudy (rsadhvika):

use compound interest formula

OpenStudy (rsadhvika):

\(A = P(1+r)^n\)

OpenStudy (rsadhvika):

A = final value P = initial value r = rate (negative here) n = number of years

OpenStudy (rsadhvika):

Value of Cheetah after 6 years :- \(A = 12000(1-.18)^6\) = ?

OpenStudy (anonymous):

C i think

OpenStudy (anonymous):

i think i put it wrong in the calculator

OpenStudy (rsadhvika):

do it again :) you need to find the Final Value for both vehicles, and see which one is greater...

OpenStudy (anonymous):

3648 cheetah

OpenStudy (rsadhvika):

Correct. What about Panther ?